an ice cube measures 2.5 cm on each side contains how many cubic centimeters of ice?
Answer by rapaljer(4671)

V=s^3
V=2.5^3 = 15.625 cubic cm.

Dr. Rapalje
Seminole State College of Florida

P.S. See my website by clicking on my tutor name "rapaljer". The link there will take you to my homepage. About halfway down my homepage, look for "Topics from Geometry," and look for the link "Perimeters, Areas, and Volumes."